SOL: SOL 3
Name of person filing report: HuntJ
- Wi-Fi Modem (lack of range). The wifi in a roughly 5 feet circle from the router, any farther and signal drops off line. We have tested this and created a signal service map by carrying laptops around the habitat to determine when it picks up signal. No known fix, will operate internet within this range for now.
- HAB Eastern Wall has sustained wind damage, panels are beginning to peel off wall. When wind blows the panels swing and hit support beam causing HAB to vibrate. In order to fix we need HAB material adhesive delivered, after which we will go EVA and patch the damage,
- ATV Yamaha 250 #3 Back Right Wheel is loose, observed to be wobbling during driving. We will assess a fix during tomorrows EVA – it is possible it simply needs to be tightened, if there is actual mechanical damage then we will move the ATV and declare it off limits.
- 3 broken old radios. All newly delivered radios are functional (thank you supply drop). Assessing broken radios to determine a fix, if none identified will mark as trash and place in a ‘broken equipment bin’. 4 EVA earpieces were declared nonfunctional, though this may have been due to incorrect settings, will continue to investigate during tomorrows EVA
Generator (hours run): 15.4 Hours
Solar— SOC % : @ 7:34 AM: 65%
@ 8:00 AM: 100 % (Generator stopped at this time)
@ 5:34 AM: 42% (Generator started at this time)
Diesel – 48% Full.
Propane – 79 Percent Volume .
Ethanol Free Gasoline (5 Gallon containers for ATV) – gallons. 2.5 Gallons. One 5 gallon tank empty, second tank half full. Request for extra fuel.
Water (trailer) – 600 Gallons
Water (static) – 500 gallons (We have water again!!)
Trailer to Static Pump used – Yes
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 35 gallons
Toilet tank emptied: No
ATV’s Used: 3 Yamaha 250’s and 1 Honda ATV used.
Oil Added? No
ATV Fuel Used: 7.5 gallons used to fill all ATVs to full, though the ATVs did not start with full tanks for our mission.
# Hours the ATVs were Used today: 02:30 hours
Notes on ATVs: See above for wobbly ATV wheel issue.
Deimos rover used: Not used
Sojourner rover used: ASSIGNED TO DIRECTOR
Spirit rover used: Not used
Opportunity rover used: Not Used
Curiosity rover used: Not Used
HabCar used and why, where? No
General notes and comments: Trailer to Static water tank was used today. Pump was dirty and needed cleaning and we did so. Actual process was nominal. We are now full on water again. When inspecting trailer tank water the front tank had a layer of FOD and detritus at the water line. We will keep watch of filter usage in case this increases our burn through rate on filters.
Summary of internet: Wifi was fixed this morning, though range is limited (see above). Please send hardware manual for the router and modem to allow us to diagnose issues and repair anything that comes up in the future (for example – to allow us to diagnose what certain light combinations signify)
Summary of suits and radios: See above
Summary of Hab operations: Brick Communication Experiment / CPR experiments successful, exercise on all crew. EVA Outer Door repaired.
Summary of GreenHab operations: No fire extinguisher observed – please deliver in future supply drop.
Summary of ScienceDome operations: Soil Salinity experiments (previous crew), dehydrating dirt for use as counterbalance and weights in exercise equipment (filling bags and medicine balls)
Summary of RAMM operations: Not Operational
Summary of health and safety issues: All nominal
Questions, concerns and requests to Mission Support:
Need extra fuel for ATVs. Of the two red fuel cans, one is empty, one is half full.
Pre-Mission food inventory is complete, will be sent as it’s own report. Please reference that report for requested supplies to address gaps in current inventory. Please give ETA for supply drop so we can adequately ration food.
Thank you for your assistance!